At 20°C, the resistance of a silver wire is 30Ω. (Coefficient of temperature=0.0038). (a) If the wire is 1.2m- long, what is the radius of the wire? (b) What is the resistance of the same wire at 10°C? (c) What is the resistivity of the wire at 25°C?




1
Expert's answer
2021-04-02T10:34:01-0400

"R=R_0(1+\alpha t)"


(a)


"R=\rho\frac{l}{A} \to A=\rho\frac{l}{R}=1.59\cdot10^{-8}\cdot\frac{1.2}{30}=636\cdot10^{-12}\ (m^2)"



"A=\pi r^2\to \frac{}{}r=\sqrt{\frac{A}{\pi}}=\sqrt{\frac{636\cdot10^{-12}}{3.14}}=1.42\cdot10^{-5}\ (m)" . Answer


(b)


"R=R_0(1+\alpha t)\to"


"R=R_0(1+0.0038\cdot20)"


"R_{10}=R_0(1+0.0038\cdot10)\to"


"R/R_{10}=(1+0.0038\cdot20)/(1+0.0038\cdot10)"


"30/R_{10}=1.0366\to R_{10}=28.94\ (\Omega)" . Answer


(c)


"\rho=\rho_0(1+\alpha t)"


"1.59\cdot10^{-8}=\rho_0(1+0.0038\cdot20)"


"\rho_{25}=\rho_0(1+0.0038\cdot25)"


"1.59\cdot10^{-8}/\rho_{25}=\rho_0(1+0.0038\cdot20)/\rho_0(1+0.0038\cdot25)\to"


"1.59\cdot10^{-8}/\rho_{25}=0.9826\to \rho_{25}=1.62\cdot10^{-8}" . Answer